Circular No. 6594

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970228 K. Hurley, Space Sciences Laboratory, Berkeley, on behalf of the Ulys-

ses Gamma-Ray Burst Team; E. Costa and M. Feroci, Istituto Astrofisica Spaziale, CNR, Frascati; F. Frontera, Universitá di Ferrara and Istituto TESRE, CNR, Bologna; D. Dal Fiume and M. Orlandini, Istituto TESRE, CNR, Bologna, report: “A refined triangulation of GRB 970228 has been done using Ulysses and SAX GRBM data. The resulting position is an annulus whose half-width is estimated to be 31′′ at 3σ confidence, or approximately one-half that previously reported by Hurley et al. (IAUC 6578). The α and δ of the annulus center and the radius are unchanged, namely α = 11^{h}03^{m}47^{s}, δ = +36^{o}31′.5 (equinox 2000.0), and 83^{o}.422. The galaxy reported by Groot et al. (IAUC 6588) lies 20′′ from the annulus center line, and is therefore consistent with this new localization.”

NOVA CASSIOPEIAE 1995 S. P. S. Eyres, Keele University; R. J. Davis and S. K. Watson, Univer-

sity of Manchester, Jodrell Bank; and M. F. Bode, Liverpool John Moores University, report the first radio detection of N Cas 1995 via MERLIN observations at 6 cm on 1996 Dec. 13 and 1997 Jan. 25: “The initial detection found an almost unresolved point with a peak of 0.48 mJy/beam, and a total flux of 0.6 mJy. By the time of the second measurement, this had risen to a peak of 0.75 mJy/beam, with a total flux of 1.2 mJy, and the nebula was marginally resolved. The peak of emission is within 0′′.18 of the Carlsberg Automatic Meridian Circle position (IAUC 6233).”

COMET C/1995 O1 (HALE-BOPP) Z. Sekanina, Jet Propulsion Laboratory, writes: “Excellent dust mod-

els have been obtained to match the diurnal evolution of the bright jet monitored by L. Jorda et al. on Feb. 28 (IAUC 6583) and the system of nearly-concentric halos observed by S. M. Larson and C. W. Hergenrother

on Feb. 8 (URL http://www.lpl.arizona.edu/bss/HB0208.html). It is

confirmed that the nucleus is not in a state of pure spin. The two features are unrelated, originating from different sources on the nucleus. Neither source is found to have been illuminated by the sun during the entire rotation cycle, both being activated only between local sunrise and sunset.”

1997 March 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green